dz Use the chain rule to find where z=(x - y)?, x=st and y=st? at
Use the Comparison Test to determine if the series is convergent or divergent. n=1 2n4-1
Answer #1

N=1 2n4-1 Now, Now, 2n-1 < 224 2n4-1 zny 2n-1 an 2n-1 an - an > bn comparison -test if diverges. Ibn diverges, then an by als® We have 2=(x-y)? 7(x-4)ox a (x-4) = 7(x-4)(1-0) az Ox a (x-41? = āx = ax 22 = 7(x-436 d (x-4)= an az 7(x-4)! (0-1) = -7(X-4